GROUND STATE X̃^{2}A′ DYNAMICS OF THE VINYL AND d_{2}-VINYL RADICALS UPON EXCITATION TO THE Ã^{2}A′′ STATE.

\underline{AARON MANN}, XIANGLING CHEN, VLADIMIR A. LOZOVSKY, C.BRADLEY MOORE, The Ohio State University, Dept. of Chemistry, 120 W. 18th Avenue, Columbus, Ohio 43210, USA.

Excitation of vinyl radical to the Ã state leads to the dissociation of vinyl radical via internal conversion into highly excited vibrational levels of the X̃ state. This was used as a method of energy deposition above the dissociation barrier of the X̃ state. Velocity maps of H and D atoms that are products of CH_{2}=CH and CD_{2}=CH dissociation accompanied by acetylene, were measured in the excitation energy range E ≈ 6750 - 9150 cm^{−1} above the dissociation barrier of the X̃ state. Rovibrational energy dependencies on the excitation energy, E, inferred for acetylene and d_{1}-acetylene products are compared with theoretically predicted ones. An upper limit of 1.3·10^{11} s^{−1} is derived for the rate constant of H-atom scrambling between the α and β carbon atoms.